Activity
From 01/21/2026 to 02/19/2026
02/16/2026
- 11:14 AM Hypernuclear ERR Task #1220 (New): Updated singles rates for detectors
- Need singles rate for occupancies and need to evaluate if we need waveform readout for some detectors in some configu...
02/12/2026
- 03:28 PM HYP Software Feature #1219 (New): Check/Generalize the tracking for single chamber
- 03:21 PM HYP Software Task #1218 (New): Kinematic reconstruction modules
- 03:21 PM HYP Software Task #1217 (New): Add Sci Fiber detector
- 11:47 AM HYP Software Task #1192 (In Progress): Update DC
- 09:28 AM HYP Software Task #1192 (In Progress): Update DC
- 11:47 AM HYP Software Task #1200 (In Progress): Work coarse processing
- 09:41 AM HYP Software Task #1200 (In Progress): Work coarse processing
- Add good hit selections and npe calculations
- 11:47 AM HYP Software Feature #1212 (In Progress): Redefine simulation input data structure
- 09:53 AM HYP Software Feature #1212 (In Progress): Redefine simulation input data structure
- Need to define and write a wrapper to pass the simulation output to fill the event object instead of CODA data.
- 10:17 AM HYP Software Feature #1210 (Resolved): Add simulation interface
- 09:50 AM HYP Software Feature #1210 (Resolved): Add simulation interface
- THcSim** classes are added to allow using simulation data as an input in the analysis chain
THcSimFile: THaRunBase... - 10:17 AM HYP Software Task #1215 (Resolved): Add new spectrometer class
- 09:56 AM HYP Software Task #1215 (Resolved): Add new spectrometer class
- Add new spectrometer class for hypernuclear spectrometers
- 09:54 AM HYP Software Support #1214 (New): Documentation
- 09:53 AM HYP Software Support #1213 (New): Test with simulation files
- 09:50 AM HYP Software Feature #1211 (New): Add digitization for new TDC modules?
- 09:48 AM HYP Software Feature #1209 (In Progress): Using simulation for replay
- 09:47 AM HYP Software Task #1208 (Resolved): Add BPM detector
- Implement BPM detector module to process FADC signal.
THcBPM is added to hcana. - 09:46 AM HYP Software Task #1207 (In Progress): Add/verify target quantity calculations
- Start with what we have in hcana
- 09:44 AM HYP Software Task #1206 (In Progress): Spectrometer
- 09:44 AM HYP Software Task #1205 (New): Review/Update DB parameters
- 09:43 AM HYP Software Task #1204 (In Progress): Reconstruction (CoarseProcessing)
- 09:43 AM HYP Software Task #1203 (Resolved): Decoding
- Implement FADC250 and vfTDC decoding
- 09:42 AM HYP Software Task #1202 (In Progress): Update TOF
- 09:41 AM HYP Software Task #1201 (New): Add track matching
- 09:40 AM HYP Software Task #1199 (Resolved): Add base class
- The idea is to have a base class to handle decoding and basic hit processing.
The reconstruction would be handled i... - 09:38 AM HYP Software Task #1198 (In Progress): Update Cherenkov
- 09:37 AM HYP Software Task #1197 (In Progress): Code clean up
- - Clean up the remaining hard-coded parts in the code
- Remove unused variables and methods - 09:36 AM HYP Software Task #1196 (In Progress): Review/Update DB parameters
- DB parameters (geometry, cuts, calibration parameters) are copied from HMS DC parameter files.
They need to be revi... - 09:35 AM HYP Software Task #1194 (Resolved): Subdector classes and reconstructed data objects
- 09:34 AM HYP Software Task #1194 (Resolved): Subdector classes and reconstructed data objects
- Added subdetector classess to handle hit processing and reconstruction at plane and chamber levels.
The modules (HY... - 09:35 AM HYP Software Task #1193 (Resolved): Add vfTDC decoder
- 09:30 AM HYP Software Task #1193 (Resolved): Add vfTDC decoder
- Support vfTDC decoding.
Start with the existing decoder used for SBS experiment. - 09:35 AM HYP Software Task #1195 (In Progress): Check ESB cosmic data
- Use ESB cosmic data for testing and debugging.
- 09:27 AM HYP Software Task #1191 (New): Hypernuclear analyzer
- Online/Offline replay software based on hcana, Podd.
The experiment uses three spectrometers (HKS, HES, ENGE). There...
02/11/2026
- Cables, Racks Layout
- Cables, Racks Layout
02/05/2026
- 11:12 AM HES Hodo Task #1158: Paddle end is fogged
- Bill prepared sandpaper(#600).
Finer sandpapers should be ordered
- 11:10 AM HES Hodo Task #1160: HES HODO's cables (HV & Signal) repariment
- Heat shrink cabling doesn't work.
Bill will ask electronics group for new connectors - 11:09 AM HES Hodo Task #1161: HES HODO Frame broken - wood branch supports top side
- Bill will solve this problem with engineering group
- 11:08 AM HKS Hodo Task #1165: Where is KTOF2Y
- KTOF2Y is low priority
02/02/2026
- 10:24 AM HES Hodo Task #1161: HES HODO Frame broken - wood branch supports top side
- Confirmed a sheared bolt is embedded in the hole.
Located the bolt head and confirmed it is made of a lightweight...
01/30/2026
- 09:53 PM HKS Hodo Task #1165 (New): Where is KTOF2Y
- KTOF1X, KTOF1Y, KTOF1Y are attached HKS detector frame, but KTOF2Y is not.
We couldn't find in ESB - 09:49 PM HKS Hodo Task #1164 (New): HKS HODO noise chcek
- PMTs to be checked if they work.
equipment : HV, oscilloscope
1. Apply high voltage
1. check noise level
H19... - 09:45 PM HKS Hodo Task #1163 (New): HKS HODO noise chcek
- PMTs to be checked if they work.
equipment : HV, oscilloscope
1. Apply high voltage
1. check noise level
H19... - 09:42 PM HKS Hodo Task #1162 (New): HKS HODO Signal Check
- HKS HODO PMT to be checked
1. Noise check
1. Gain check w/ beta ray
1. Time resolution measurement - 09:40 PM HES Hodo Task #1161 (New): HES HODO Frame broken - wood branch supports top side
- One of HES HODO frame bolts are missing (HES ch30-58 frame's TOP side)
Temporary, wood branch supports instead of th... - 09:29 PM HES Hodo Task #1160: HES HODO's cables (HV & Signal) repariment
- Check which cables to be repaired
|_.Ch# T(TOP) or B(Bottom)|_.12B|_.39B|_.43B|_.44T|_.46B|_.48T|_.48B|_.54T|_.56T|_... - 09:17 PM HES Hodo Task #1160 (New): HES HODO's cables (HV & Signal) repariment
- Some HV cables and Signal cables are uncovered, especially on the connection of PMT or port.
Additional covers are n...
01/29/2026
- 12:02 PM HES Hodo Task #1159 (New): Soft Black Sheet
- JLab's black sheet is paper-like, hard one.
We prefer softer one which we conventionally used in Japan.
Someone sho... - 11:58 AM HES Hodo Task #1158 (New): Paddle end is fogged
- Some paddles has fogged end.
We have to decide how to clean - 11:44 AM HES Hodo Task #1123: Connector Reinforce
- https://redmine.jlab.org/issues/1157
- 11:38 AM HES Hodo Task #1157 (New): HES HODO Repairment : PMT-Light Guide-Paddle connection
- HES HODO Paddles have to be repaired.
PMT - Light Guide or Light Guide - Paddle connections are broken
!%E3%82%B9... - 11:16 AM HES Hodo Task #1156: HES HODO frame parts are missed
- https://redmine.jlab.org/issues/1155
- 11:13 AM HES Hodo Task #1156: HES HODO frame parts are missed
- |_.Seg #|_.1|_.2|_.3|_.8|_.29|_.30|_.46|_.51|_.54|_.56|._58|_.Total |
|TOP Black Front | 1|1 | |1 |1 |1 |1 |1 ... - 10:53 AM HES Hodo Task #1156 (New): HES HODO frame parts are missed
- HES HODOs are used for trigger counters in ESB.
Metal frame parts (black and silver) fixing HODOs on HES large fram...
01/27/2026
- 04:19 PM Task #1155 (New): Hodoscope commissioning with cosmic ray in ESB
- Commissioning of hodocsopes, HKS, HES, and ENGE, is onging in ESB.
HES Hodoscope: https://docs.google.com/document/... - 11:16 AM Task #1154 (New): Drift chamber commissioning with cosmic ray ESB
- Commissioning of drift chambers, EDC and KDC, is onging in ESB.
EDC ELOG: https://docs.google.com/document/d/1SCD2p...
Also available in: Atom